In accounting:

  1. What is earnings per share (EPS)? What does it measure?  
  2. Is it better to have a higher or lower EPS? Why?
  3. Is it better to have a higher or lower price-earnings ratio (P-E ratio)? State your reasoning.
  4. Is a low PE ratio good and why? What would be considered a high PE ratio
Answer #1

1) EPS - Earnings available to each equity share

Formula of EPS
EPS = Earnings available to shareholders/No of shares

2) No. lower EPS represents lower earnings per share (Lower returns for investment)

3 ) Lower P-E ratio is better.
P.E ratio = Market price/EPS
P-E means how many times share price trading in market compared to EPS
Lower P-E gives more return

4) Lower P-E gives more return on investment
Example: Case -1: EPS = $2, P-E ratio = 10 times, MPS = $2*10 = $20
In this case investor get $2 return for $20 investment (10% return on investment)

Case -2: EPS = $2, P-E ratio = 5 times, MPS = $2*10 = $10
In this case investor get $2 return for $10 investment (20% return on investment)

In case 2 investor will get 20% return when PE is only 5 times compared to 10% return when PE is 5 times

Therefore when PE ratio is low return will be high